2x over y minus 4z
X=9 y=3 z= -1

Found 2 solutions by persian52, user_dude2008:
Answer by persian52(161)

(2(9))/(3)-4(-1)
Multiply 2 by 9 in the numerator.
(2*9)/(3)-4(-1)
Multiply 2 by 9 to get 18.
(18)/(3)-4(-1)
Reduce the expression (18)/(3) by removing a factor of 3 from the numerator and denominator.
6-4(-1)
Multiply -4 by each term inside the parentheses.
6+4
Add 4 to 6 to get 10.
Answer: 10

Answer by user_dude2008(1862)

(2x)/(y-4z)=(2*9)/(3-4(-1)) = 18/(3+4) = 18/7

Answer: (2x)/(y-4z)=18/7